# references.bib

### v1

@@ 6. Discussion
We have found a few exceptions to this trend. Firstly, CacheSlicer (shahrad_provisioning_2021) is a proof-of-concept of an LLC-aware extension of the Microsoft Azure virtual machine scheduler. Secondly, Zeus (zhang_zeus_2021) attempts to improve upon Kubernetes resource management and increase cluster utilization at NetEase. Finally, the intent-driven orchestrator (metsch_intent-driven_2023) tries to learn how changes in resource allocations affect service metrics and then make appropriate adjustments.
@@ 6.1. Privacy Concerns of Private and Public Clouds
The type of cloud operated has a big impact on the type of approaches usable in a given scenario. A private cloud, hosted on-premise, running only software written in-house has vastly different requirements than a cloud running customer software. In the first scenario, cloud operators are able to monitor applications in any way imaginable. This includes prior-to-deployment offline profiling of software. This allows resource management controllers with full application-aware integration to be employed, as leveraged by the intent-driven orchestrator (metsch_intent-driven_2023). However, companies hosting other organizations’ software through infrastructure-as-a-service (IaaS) solutions are required to take additional security measures into account as detailed by shahrad_provisioning_2021 (shahrad_provisioning_2021). In such scenarios, as in the context of public cloud providers, there is no alternative but to treat customer software as black boxes. Nonetheless, managed platform-as-a-service (PaaS) offerings could provide software add-ons allowing customers to provide a resource manager with application-level metrics and targets. Such solutions would make customer software a little more “grey” without becoming an intellectual property protection concern.
Additionally, there may be sensitive data managed by the application that must be handled with discretion due to legal compliance reasons, such as with the EU’s General Data Protection Regulation (GDPR) (noauthor_general_nodate).
@@ 6.2. Consolidation of Workloads
Many of the workload consolidation approaches discuss consolidation from the point of view that workloads are either latency-critical (LC) continuously running services or best-effort (BE) jobs. The general idea is to ensure SLOs are met by one (lo_heracles_2015; zhao_rhythm_2020) or more (chen_parties_2019; patel_clite_2020; nikas_dicer_2019; qin_locpart_2017; qiu_slo_2021; li_rldrm_2020; zhu_dirigent_2016; chen_avalon_2019) LC services while utilizing additional machine resources to pack and maximize the performance of BE jobs. This way, the utilization of hardware can be increased while performance expectations of services are met. While the way different types of workloads are used to improve system utilization is sound, the approach assumes a decently static amount of LC traffic and a steady stream of BE jobs awaiting processing. They do not consider scenarios where LC traffic may vary enough that replication scaling is preferred or required, as is common in cloud-based environments. Nor is there any consideration for the underprovisioning that would emerge if no BE jobs are scheduled.
@@ 6.3. Classification of Workloads
As may be observed by tables 4 and 5, many publications include some form of workload classification as part of their proposed solutions. This is a reasonable way of tackling the resource partitioning problem due to the constraint of finite classes of service (CLOS) available in hardware and subsequently the resctrl resource management interface. Therefore, workloads are required to be generically grouped together according to some scheme. This may range from a simple binary distinction between resource-sensitive and insensitive (e.g. CA (pons_improving_2018)) to more sophisticated program phase classification schemes (e.g. dCat (xu_dcat_2018)).
@@ 6.4. QoS Enforcement Feature Strengths and Weaknesses
Given the numerous studies that have successfully applied both cache allocation and memory bandwidth limitation to achieve their respective goals (be it performance, fairness, energy efficiency, utilization, or a combination of the aforementioned goals), it is clear that the hardware QoS enforcement features surveyed in this article are appropriate tools to utilize when optimizing for such goals. Additionally, the aforementioned features have also been shown to be effective in mitigating other issues such as security threats (liu_catalyst_2016).
@@ 7. Research Gap Analysis
As we discussed in Section 6, it is highly likely that public cloud providers make extensive use of the QoS enforcement features found in recent server processors. However, our survey of the academic landscape has made it clear that there is space for holistic solutions that are able to intelligently manage cache and memory bandwidth in the context of multiple machines or even entire data centers. Only a few publications approach QoS management in the context of data centers (shahrad_provisioning_2021; jeatsa_casy_2022), which focus solely on cache partitioning. This leaves room for additional research that could improve the utilization of hardware in smaller, on-premises or edge deployments that generally cannot benefit from the advancements made behind public cloud providers’ closed doors. A fitting approach would be to extend service orchestrators such as Kubernetes (the_kubernetes_authors_kubernetes_2022) with intelligent cache and memory bandwidth awareness.
Observing the evaluations listed in Table 3, it is notable that no evaluation focuses on the specifics of AMD’s QoSE implementation of hardware QoS enforcement features. This is significant as studies have shown inconsistent behavior between hardware generations (sohal_closer_2022), even from the same manufacturer (Intel). Thus, a systematic evaluation where the behavior of AMD’s implementation should be conducted. This would be nicely followed by a comparison of the behaviors of the two chip manufacturers’ implementations.
@@
Furthermore, by assuming that most LC cloud workloads are allocated more resources than required to account for sudden spikes in load without breaking SLOs, an opportunity to improve system-wide efficiency by utilizing this gap between allocation and actual utilization may be observed. Limiting workloads that operate with some margin to their SLOs with intelligent management of shared resources should allow net efficiency gains to be made without impacting the end-user experience. This concept is in part addressed by pang_adaptive_2021 (pang_adaptive_2021). However, their approach only rudimentarily utilizes cache partitioning without touching memory bandwidth limiting.
@@ 8. Concluding Remarks
We have conducted a broad survey of the research landscape and identified how academics are utilizing hardware-based quality-of-service enforcement features implemented in commodity x86 processors. Through our use of a reference snowballing approach, a search space consisting of $3685$ references was reviewed. Out of which, $79$ papers were identified to meet our specified inclusion criteria. We proposed a novel taxonomy of the included papers that structures and classifies the current approaches used by scholars. By observation of this taxonomy, conclusions could be drawn regarding the current general directions of research. Furthermore, gaps in current state-of-the-art research were identified and detailed.
@@ Acknowledgements.
The authors would like to thank Ravi Iyer (Intel Labs) for his input and invaluable feedback. Financial support has been provided in part by the Knut and Alice Wallenberg Foundation under grant KAW 2019.0352 and by the eSSENCE Programme under the Swedish Government’s Strategic Research Initiative.
[…]

### v2

@@ 6. Discussion
We have found a few exceptions to this trend. Firstly, CacheSlicer (shahrad_provisioning_2021, 93) is a proof-of-concept of an LLC-aware extension of the Microsoft Azure virtual machine scheduler. Secondly, Zeus (zhang_zeus_2021, 121) attempts to improve upon Kubernetes resource management and increase cluster utilization at NetEase. Finally, the intent-driven orchestrator (metsch_intent-driven_2023, 64) tries to learn how changes in resource allocations affect service metrics and then make appropriate adjustments.
@@ 6.1. Privacy Concerns of Private and Public Clouds
The type of cloud operated has a big impact on the type of approaches usable in a given scenario. A private cloud, hosted on-premise, running only software written in-house has vastly different requirements than a cloud running customer software. In the first scenario, cloud operators are able to monitor applications in any way imaginable. This includes prior-to-deployment offline profiling of software. This allows resource management controllers with full application-aware integration to be employed, as leveraged by the intent-driven orchestrator (metsch_intent-driven_2023, 64). However, companies hosting other organizations’ software through infrastructure-as-a-service (IaaS) solutions are required to take additional security measures into account as detailed by (shahrad_provisioning_2021, 93) (shahrad_provisioning_2021, 93). In such scenarios, as in the context of public cloud providers, there is no alternative but to treat customer software as black boxes. Nonetheless, managed platform-as-a-service (PaaS) offerings could provide software add-ons allowing customers to provide a resource manager with application-level metrics and targets. Such solutions would make customer software a little more “grey” without becoming an intellectual property protection concern.
Additionally, there may be sensitive data managed by the application that must be handled with discretion due to legal compliance reasons, such as with the EU’s General Data Protection Regulation (GDPR) (noauthor_general_nodate, 82).
@@ 6.2. Consolidation of Workloads
Many of the workload consolidation approaches discuss consolidation from the point of view that workloads are either latency-critical (LC) continuously running services or best-effort (BE) jobs. The general idea is to ensure SLOs are met by one (lo_heracles_2015, 61, 123) or more (chen_parties_2019, 17, 76, 70, 83, 85, 57, 125, 14) LC services while utilizing additional machine resources to pack and maximize the performance of BE jobs. This way, the utilization of hardware can be increased while performance expectations of services are met. While the way different types of workloads are used to improve system utilization is sound, the approach assumes a decently static amount of LC traffic and a steady stream of BE jobs awaiting processing. They do not consider scenarios where LC traffic may vary enough that replication scaling is preferred or required, as is common in cloud-based environments. Nor is there any consideration for the underprovisioning that would emerge if no BE jobs are scheduled.
@@ 6.3. Classification of Workloads
As may be observed by tables 4 and 5, many publications include some form of workload classification as part of their proposed solutions. This is a reasonable way of tackling the resource partitioning problem due to the constraint of finite classes of service (CLOS) available in hardware and subsequently the resctrl resource management interface. Therefore, workloads are required to be generically grouped together according to some scheme. This may range from a simple binary distinction between resource-sensitive and insensitive (e.g. CA (pons_improving_2018, 80)) to more sophisticated program phase classification schemes (e.g. dCat (xu_dcat_2018, 111)).
@@ 6.4. QoS Enforcement Feature Strengths and Weaknesses
Given the numerous studies that have successfully applied both cache allocation and memory bandwidth limitation to achieve their respective goals (be it performance, fairness, energy efficiency, utilization, or a combination of the aforementioned goals), it is clear that the hardware QoS enforcement features surveyed in this article are appropriate tools to utilize when optimizing for such goals. Additionally, the aforementioned features have also been shown to be effective in mitigating other issues such as security threats (liu_catalyst_2016, 60).
@@ 7. Research Gap Analysis
As we discussed in Section 6, it is highly likely that public cloud providers make extensive use of the QoS enforcement features found in recent server processors. However, our survey of the academic landscape has made it clear that there is space for holistic solutions that are able to intelligently manage cache and memory bandwidth in the context of multiple machines or even entire data centers. Only a few publications approach QoS management in the context of data centers (shahrad_provisioning_2021, 93, 51), which focus solely on cache partitioning. This leaves room for additional research that could improve the utilization of hardware in smaller, on-premises or edge deployments that generally cannot benefit from the advancements made behind public cloud providers’ closed doors. A fitting approach would be to extend service orchestrators such as Kubernetes (the_kubernetes_authors_kubernetes_2022, 100) with intelligent cache and memory bandwidth awareness.
Observing the evaluations listed in Table 3, it is notable that no evaluation focuses on the specifics of AMD’s QoSE implementation of hardware QoS enforcement features. This is significant as studies have shown inconsistent behavior between hardware generations (sohal_closer_2022, 95), even from the same manufacturer (Intel). Thus, a systematic evaluation where the behavior of AMD’s implementation should be conducted. This would be nicely followed by a comparison of the behaviors of the two chip manufacturers’ implementations.
@@
Furthermore, by assuming that most LC cloud workloads are allocated more resources than required to account for sudden spikes in load without breaking SLOs, an opportunity to improve system-wide efficiency by utilizing this gap between allocation and actual utilization may be observed. A traditional approach to this problem is to use overcommitment (cohen_overcommitment_2019, 21), but this can lead to unpredictable performance degradation due to increased resource congestion. A possible hardware-level solution could be to limit workloads that operate with some margin to their SLOs with intelligent management of QoS enforcement, which should allow net efficiency gains to be made without impacting the end-user experience. This concept is in part addressed by (pang_adaptive_2021, 72) (pang_adaptive_2021, 72) which indicates the feasibility of the technique. However, their approach only rudimentarily utilizes cache partitioning without touching memory bandwidth limiting.
@@ 8. Concluding Remarks
We have conducted a broad survey of the research landscape and identified how academics are utilizing hardware-based quality-of-service enforcement features implemented in commodity x86 processors. Through our use of a reference snowballing approach, a search space consisting of $3685$ references was reviewed. Out of which, $78$ papers were identified to meet our specified inclusion criteria. We proposed a novel taxonomy of the included papers that structures and classifies the current approaches used by scholars. By observation of this taxonomy, conclusions could be drawn regarding the current general directions of research. Furthermore, gaps in current state-of-the-art research were identified and detailed.
@@ Acknowledgements.
We would like to thank our anonymous reviewers as well as Ravi Iyer for their input and invaluable feedback. Financial support has been provided in part by the Knut and Alice Wallenberg Foundation under grant KAW 2019.0352 and by the eSSENCE Programme under the Swedish Government’s Strategic Research Initiative.